OpenStudy (anonymous):

(9^2y)/3^(7-y) = (3^(4y+3))/27^(y-2) ??? y = ??? I've got 8/3.

OpenStudy (anonymous):

the answer is y=4

OpenStudy (anonymous):

how did you get that??

OpenStudy (anonymous):

on the LHS we have 3^4y/3^7-y which i s 3^5y-7 on the rhs we have 3^4y+3/3^3y-6 which is 3^y+9 equating the powers on the 2 sides we get y=4

jhonyy9 (jhonyy9):

like expoent will get 5y-7=y+9 from this you need calcule the y so 4y=16 so y=16/4 =4 y=4

jhonyy9 (jhonyy9):

like exponent wann to be - sorry

OpenStudy (anonymous):

got it already?

jhonyy9 (jhonyy9):

understandably ?

jhonyy9 (jhonyy9):

tanvidais13 - ok ?

jhonyy9 (jhonyy9):

so 8/3 not is right

OpenStudy (anonymous):

ahhhhhh, thank you SOOO much :) when i was making the base of 9 to 3, while doing the (3^2)^2y, i accidentally got 3^6y :P

jhonyy9 (jhonyy9):

ok good luck bye

OpenStudy (anonymous):

thank youuu, and same you to as well :D

OpenStudy (anonymous):

it took great deal of time

OpenStudy (anonymous):

really?? It took me only a little time, but I DID get it wrong, so.... yeahh, i'm weird :P

OpenStudy (anonymous):

haha and funny too